# Characterization of birefringence inhomogeneity of KAGRA sapphire mirrors from transmitted wavefront error measurements

### v1

@@ Appendix A Effective Jones matrix of a non-uniform medium along the optical axis
In Section II, we assumed that the transmission of light through a substrate can be described by a single Jones matrix (Equation 2). At a glance, this seems to imply that we also assume the birefringence properties ($\theta$, $n_{e}$ and $n_{o}$) to be constant along the depth of the substrate. However, this assumption does not hold for most real substrates. For the sapphire substrates of the KAGRA mirrors, for example, the fluctuation of birefringence is supposed to come from local mis-alignment of crystal axes, possibly caused by residual internal stress, lattice defects and so on. Then, the birefringence properties must vary not only laterally but also along the depth of the substrate. Then a question arises: can we use a single Jones matrix to describe a complex birefringent medium? In this section, we will address this question.
@@
We start from splitting a thick substrate into thin slices as shown in Figure 11. Each slice is thin enough that the birefringence properties can be regarded as constant within it. Transmission of light through such a thin slice can be described by a single Jones matrix of the form in Equation 2.
As we make the slice infinitesimally thin, $\alpha_{-}$ becomes also infinitesimally small. From Equation 2, we can expand the Jones matrix with $\alpha_{-}$ and leave only its first order terms (the common phase term $\alpha_{+}$ is omitted here),
@@
The next step is to consider the transmission of light through two consecutive slices having different Jones matrices:
@@
By comparing Equation 39 and Equation 30, we need to find $\alpha_{\mathrm{e}}$ and $\theta_{\mathrm{e}}$ which satisfy the following conditions,
@@
Finding such a combination of $\alpha_{\mathrm{e}}$ and $\theta_{\mathrm{e}}$ is always possible. Then we can express $M_{12}$ using these parameters as,
@@
This is an effective Jones matrix for the transmission of two consecutive thin slices. By repeating this process, we can always find a single Jones matrix of the form shown in Equation 2 representing the transmission of a thicker substrate. Therefore, the use of Equation 2 for the analysis of an inhomogeneous thick substrate is justified.
@@ Appendix B Influence of tilt of TWE maps on the estimation of beam shape and beam power
In Section III.2, we show that the piston, which is usually ignored in an interferometric measurement with Fizeau interferometers, however, is an important factor in birefringence measurements with the method described in this paper. The tilt, also ignored in TWE measurements, has a similar effect. In this section, we give an example of its influence to our birefringence characterizations.
The tilt of a surface can be described by a combination of the first-order Zernike polynomials $Z_{1}^{-1}$ and $Z_{1}^{1}$ (shwon in Figure 12). Thus we need to add two degrees of freedom of tilt for nominator and denominator respectively in Equation 23 and this ends up with four more degrees of freedom ($a$, $b$, $c$ and $d$) for parameter scan, which yields
@@
Figure 12: The first order coefficients of Zernike polynomials. Any tilt of a surface can be described with a combination of them with different amplitudes.
According to Figure 6, the change in the nominator in Equation 43 mainly influences the continuity factor, while the change in the denominator has a larger effect on the beam shape and power. Here, we set $a=b=0$ to see the influence of tilt on beam shape and power. Pistons $p_{1}$ and $p_{2}$ are fixed according to values giving results in red squares in Figure 6. Figure 13 shows the change of beam shape and power when we apply different amount of tilt to the denominator. The results labeled in red squares show a good match with birefringence measurements at POP in terms of both beam shape and beam power in p-polarization.
Figure 13: An example of parameter scan showing the change of beam shape and power in p-polarization with different magnitude of tilt.

### v2

@@ Appendix A Effective Jones matrix of a non-uniform medium along the optical axis
In Section II, we assumed that the transmission of light through a substrate can be described by a single Jones matrix (Equation (2)). This seems to imply that we also assume the birefringence properties ($\theta$, $n_{e}$ and $n_{o}$) to be constant along the depth of the substrate. However, this assumption does not hold for most real substrates. For the sapphire substrates of KAGRA mirrors, for example, the fluctuation of birefringence is supposed to originate from the local misalignment of crystal axes, possibly caused by residual internal stress and lattice defects. Therefore, the birefringence properties must vary not only laterally but also along the depth of the substrate. Thus, the question arises: can we use a single Jones matrix to describe a complex birefringent medium? This section addresses this question.
@@
We began by splitting a thick substrate into thin slices, as shown in Figure 11. Each slice is sufficiently thin such that the birefringence properties can be regarded as constant. The transmission of light through such a thin slice can be described by a single Jones matrix, as shown in Equation (2).
As we make the slice infinitesimally thin, $\alpha_{-}$ becomes also infinitesimally small. From Equation (2), we can expand the Jones matrix with $\alpha_{-}$ and leave only its first order terms (the common phase term $\alpha_{+}$ is omitted here),
@@
Next, we considered the transmission of light through two consecutive slices with different Jones matrices:
@@
By comparing Equation (29) and Equation (26), we find $\alpha_{\mathrm{e}}$ and $\theta_{\mathrm{e}}$, which satisfy the following conditions:
@@
Finding such a combination of $\alpha_{\mathrm{e}}$ and $\theta_{\mathrm{e}}$ is always possible. Then we express $M_{12}$ using these parameters as follows:
@@
This is an effective Jones matrix for transmitting two consecutive thin slices. By repeating this process, we can obtain a single Jones matrix of the form shown in Equation (2), which represents the transmission of a thick substrate. Therefore, the use of Equation (2) for the analysis of thick, inhomogeneous substrates is justified.
@@ Appendix B Influence of tilt of TWE maps on the estimation of beam shape and beam power
Figure 12: The first order coefficients of Zernike polynomials. Tilts of a surface can be described using a combination of them with different amplitudes.
Figure 13: An example of parameter scan showing the change of beam shape and power in p-polarisation with different magnitude of tilt.
In Section III.2, we show that the piston, which is typically ignored in interferometric measurements using Fizeau interferometers, is an important factor in birefringence measurements using the method described in this paper. Tilt, which is also ignored in TWE measurements, has a similar effect. In this section, we provide an example of its influence on birefringence characterisation.
The tilt of the surface can be described by a combination of the first-order Zernike polynomials $Z_{1}^{-1}$ and $Z_{1}^{1}$ (shwon in Figure 12). Therefore, we need to add two degrees of freedom of tilt for the numerator and denominator in Equation (23), which results in four more degrees of freedom ($a$, $b$, $c$ and $d$) for the parameter scan, yielding
@@
According to Figure 6, the change in the numerator of Equation (33) mainly influences the continuity factor, whereas the change in the denominator has a larger effect on the beam shape and power. Here, we set $a=b=0$ to determine the influence of the tilt on the beam shape and power. Pistons $p_{1}$ and $p_{2}$ are fixed according to their values, resulting in the red squares shown in Figure 6. Figure 13 shows the change in the beam shape and power when different amounts of tilt are applied to the denominator. The results labelled with red squares show a good match with the birefringence measurements at the POP in terms of both the beam shape and beam power in p-polarisation.